# Beyond correlation: The impact of human uncertainty in measuring the effectiveness of automatic evaluation and LLM-as-a-judge

### v2

@@ 3 Analysis Settings and Results
The binned JSD for ordinal perception data has 2 main advantages, (1) it measures how closely machine generated labels mimic human perception without the need for a single gold label (2) human and machine labels are not treated interchangeably, as the items in a given bin are selected by the human median or majority value. For HM correlation, the proposed approach can be applied at item-level when the number of labels collected per item is high enough to form a distribution, however, for judging model generated output even obtaining 3 human labels per item can be expensive. Hence, we suggest computing the distribution distance per bin. The main shortcoming of this approach is that some key item level discrepancies between machines and humans ratings might not be clearly surfaced. For instance, for a given item, if all the humans assign a rating of 1 and machines assign say 4, at an aggregate level this pattern may not be obvious if this pattern is not frequent. Hence, we also recommend reporting correlations stratified by uncertainty in human labels, e.g., for items that have high agreement labels, such as perfect human agreement as shown in Table 2.
@@ 4.2 Challenges in metrics and interpretability
In addition, statistical analysis, such as null-hypothesis and significance testing, is essential for determining whether one model outperforms another by random chance. Here, the chance component includes 2 aspects, (1) chance due to the nature of samples in the evaluation set (2) uncertainty in human labels. A third aspect, even harder, is estimating the error rate as a result of systematically unpredictable erroneous labels from any automated evaluator. Future studies should explore these problems, including approaches like resampling (Deutsch et al., 2021). Incorporation of chance in rank correlation is also an important aspect to account for when two models differ in rank, but the corresponding difference in absolute scores is negligible, then the difference in the rank may not be meaningful.
@@ A.7 Code for Binned-JSD-perception
Human labels Human median (Bin) Model predictions
2, 2, 3 2 3, 2
1, 2, 2 2 1, 1
2, 3, 3 3 2, 2
@@
A.8 Visualization of preference data
@@
A.9 Extended table results
@@
A.10 Random Simulation Code
A.10.1 Random Simulation Code - Binary classification (Nominal data)
@@
A.10.2 Random Simulation Code - 3 way classification (Ordinal data)
@@ A.11 Case study
Topical chat (Mehri & Eskenazi, 2020) 60 Uses Knowledge 3 (0, 1)
@@
Table 7: Datasets summary
@@ A.11.1 Analysis of results
    Impact of stratification: When we compare the overall performance (column All in Table 8) on dataset TC (criteria understandable) with dataset QAGS, Randolph-$\kappa$ drops substantially by 19 points (0.93 $\rightarrow$ 0.74). However, QAGS dataset has around 66% of the samples that have perfect human agreement, while TC has 96% of the sample with human agreement. When we compare the samples with perfect human agreement between the 2 datasets, the model performance gap reduces to just 6 points (0.93 $\rightarrow$ 0.87). With DICES dataset (crowdsourced with over 100 annotations per item, with no perfect agreement items), on the other hand, the model seems to struggle across all metrics and stratification groups, indicating much deeper investigation is required. The general trend of where in a stratified group with higher proportion of noisy or uncertain samples (as measured by low HH correlation), the HM correlation seems to outperform HH correlation as highlighted in Table 8 applies as previously discussed in Section 3.1.
A.12 Toy example: Application of binned-JSD
Human labels Human median ($\overline{H}$) Good model Poor model
2, 2, 3 2 3 1
1, 2, 2 2 1 3
2, 3, 3 3 2 4
Table 9: A toy dataset to illustrate the use of binned-jsd
Metric Good Model Poor Model Does better model score better
Kendall-$\tau$ 0.00 0.82 False
Spearman-$\rho$ 0.00 0.87 False
Krippendorff-$\alpha$0.06 -0.06 False
$JS_{b}$ 0.56 0.65 True

### v3

@@ 3 Analysis Settings and Results
The binned JSD for ordinal perception data has 2 main advantages, (1) it measures how closely machine labels mimic human perception without the need for a single gold label (2) human and machine labels are not treated interchangeably, as the items in a given bin are selected by the human median or majority value. For HM correlation, the proposed approach can be applied at item-level when the number of labels collected per item is high enough to form a distribution, however, for judging model generated output even obtaining 3 human labels per item can be expensive. Hence, we suggest computing the distribution distance per bin. The main shortcoming of this approach is that some key item level discrepancies between machines and humans ratings might not be clearly surfaced. For instance, for a given item, if all the humans assign a rating of 1 and machines assign say 4, at an aggregate level this pattern may not be obvious if this pattern is not frequent. Hence, we also recommend reporting correlations stratified by uncertainty in human labels, e.g., for items that have high agreement labels, such as perfect human agreement as shown in Table 2.
@@ 4.2 Challenges in metrics and interpretability
In addition, statistical analysis, such as null-hypothesis and significance testing, is essential for determining whether one model outperforms another by random chance. Here, the chance component includes 2 obvious aspects, (1) chance due to the nature of samples in the evaluation set (2) uncertainty in human labels. A third aspect, even harder, is estimating the error rate as a result of systematically unpredictable erroneous labels from any automated evaluator. Future studies should explore these problems, including approaches like resampling (Deutsch et al., 2021). Incorporation of chance in rank correlation is also an important aspect to account for when two models differ in rank, but the corresponding difference in absolute scores is negligible, then the difference in the rank may not be meaningful.
@@ A.7 Code for Binned-JSD-perception
Id Human labels Human median (Bin) Model predictions
A 2, 2, 3 2 3, 2
B 1, 2, 2 2 1, 1
C 2, 3, 3 3 2, 2
@@
A.8 Toy example: Application of binned-JSD and where it can a better metric
Human labels Human median ($\overline{H}$) Good model Poor model
2, 2, 3 2 3 1
1, 2, 2 2 1 3
2, 3, 3 3 2 4
Table 5: A toy dataset to illustrate the use of binned-jsd
Metric Good Model Poor Model Does better model score better
Kendall-$\tau$ 0.00 0.82 False
Spearman-$\rho$ 0.00 0.87 False
Krippendorff-$\alpha$0.06 -0.06 False
$JS_{b}$ 0.56 0.65 True
A.9 Visualization of preference data
@@
A.10 Extended table results
@@
A.11 Random Simulation Code
A.11.1 Random Simulation Code - Binary classification (Nominal data)
@@
A.11.2 Random Simulation Code - 3 way classification (Ordinal data)
@@ A.12 Case study
60 Uses Knowledge 3 (0, 1)
@@
Table 9: Categorical datasets summary
@@ A.12.1 Analysis of results
    Impact of stratification: When we compare the overall performance (column All in Table 10) on dataset TC (criteria understandable) with dataset QAGS, Randolph-$\kappa$ drops substantially by 19 points (0.93 $\rightarrow$ 0.74). However, QAGS dataset has around 66% of the samples that have perfect human agreement, while TC has 96% of the sample with human agreement. When we compare the samples with perfect human agreement between the 2 datasets, the model performance gap reduces to just 6 points (0.93 $\rightarrow$ 0.87), pointing to how comparing datasets with different proportion of uncertain samples can affect our conclusion (in this case incorrectly that the model performance is substantially lower in QAGS compared to TC). With DICES dataset (crowdsourced with over 100 annotations per item, with no perfect agreement items), on the other hand, the model seems to struggle across all metrics and stratification groups, indicating much deeper investigation is required. The general trend of where in a stratified group with relatively higher proportion of uncertain samples (as measured by low HH correlation), the $\text{H}^{w}\text{M}^{w}$ correlation seems to outperform HH correlation as highlighted, as shown in Table 10 also applies, indicating to how models can superficially appear to approximate human majority when proportion samples of uncertainty is relatively higher, demonstrating the impact of RQ1 discussed in section 3.1.
A.12.2 Analysis of ordinal data
Firstly, looking at Table 11, superficially looking at the aggregate scores ( Row marked “All” across all 4 criteria) it might appear as though (a) absolute human to machine correlation is low (columns $\overline{\text{H}}\overline{\text{M}}$) (b) when we compare HM correlation to HH, $\overline{\text{H}}\overline{\text{M}}$ is similar to HH correlation.
Dataset Size Criteria # Annotations Choices
Topical chat (Mehri & Eskenazi, 2020) 60 engaging 3 1,2,3
60 maintains context 3 1,2,3
60 natural 3 1,2,3
60 overall 3 1,2,3,4,5
Krippendorff’s-$\alpha$ $JS_{b}$ $\rho$ $\tau$
C Partition Size (%) HH MM $\overline{\text{H}}\overline{\text{M}}$ $\Delta$ HM $\overline{\text{H}}\overline{\text{M}}$ $\overline{\text{H}}\overline{\text{M}}$
ENG All 100.00 0.03 0.72 -0.02 0.05 0.24 0.25 0.24
PA=1 75.00 1.00 0.66 -0.14 1.14 0.34 - -
60% $\leq$ PA $<$ 80% 25.00 -0.39 0.82 -0.01 -0.38 0.15 0.26 0.25
MC All 100.00 -0.06 0.74 -0.07 0.01 0.13 - -
PA=1 81.67 1.00 0.65 -0.04 1.04 0.20 - -
60% $\leq$ PA $<$ 80% 18.33 -0.45 0.78 -0.17 -0.28 0.07 - -
NAT All 100.00 -0.07 0.71 -0.20 0.13 0.24 - -
PA=1 78.33 1.00 0.75 -0.18 1.18 0.35 - -
60% $\leq$ PA $<$ 80% 21.67 -0.46 0.59 -0.25 -0.21 0.14 - -
OV All 100.00 0.10 0.75 -0.28 0.38 0.33 -0.17 -0.16
PA=1 61.67 1.00 0.76 -0.21 1.21 0.41 - -
60% $\leq$ PA $<$ 80% $3$0.31 0.72 -0.48 0.17 0.35 -0.37 -0.35
0 $\leq$ PA $<$ 40% 5.00 -0.32 0.66 -0.25 -0.07 0.24 - -
Table 11: Analysis of performance on the Topical chat dataset on criteria Engaging (ENG), Maintains context (MC), Naturalness (NAT) and Overall (OV). Overall, all the metrics that measure human – to machine correlation, across all metrics, seem quite low. At this point, it is quite difficult to comprehend what is causing this, until we visualize them as shown in Figures 8 – Figures11
In order to investigate the poor $\overline{\text{H}}\overline{\text{M}}$, we perform the following analysis.
  1. 1.
    Is the $\overline{\text{H}}\overline{\text{M}}$ low because the model predictions are indeed quite different to humans? Investigating the criteria Engaging, as shown in Table 11, the $\overline{\text{H}}\overline{\text{M}}$ correlation is low (Krippendorff-$\alpha$0.02, $\rho$ 0.25, and $\tau$ 0.24). When we visualize this data, we find that over 96% of the human labels have a median of 3, and only 3.3% (2 out of 60) have a different median value ($\overline{\text{H}}=2$), as shown in Figure 8. Obviously, there is a considerable class imbalance, that is likely to lower Krippendorff-$\alpha$. In the case of rank correlation, over 96% of $\overline{\text{H}}=3$, and of these roughly 30% have been assigned a score of 3 $\overline{\text{M}}=3$, lowering rank correlation metrics.
    Given this imbalance, there are 2 scenarios – (a) human annotation is reliable and the LLM prompts need to be optimized to improve the performance (b) Human annotation guidelines need to be revised. If we assume that the human annotation is reliable, then the actionable step here is to ensure that the LLM prompt is optimized such that the items are incorrectly labelled as 2, are correct to 3.
    We can see a similar pattern in criteria MC (see Figure 9), NAT (see Figure 10). For criteria OV, in LJ labels are quite different to machine labels (see Figure 11).
Figure 8: Topical chat criteria – Engaging: Over 96% of the samples are assigned a median score of 3 by humans fairly consistently, where around 90% of the human label bin $\overline{H}=3$ have the same value as the median. Only 2 (3.3% of 60) items are in bin $\overline{H}=2$. There are no human labels for ordinal value 1, as seen in bins 1, 2 and 3. This shows how unbalanced the human labels are.
Figure 9: Topical chat criteria – Maintains context. The human median is 3 for 100% of the samples. Hence, rank correlation is not applicable . While in majority of the case (80% of the samples), human median seems to match the machine median, k-$\alpha$ correlation $\overline{\text{H}}\overline{\text{M}}$ is quite low at -0.07 as shown Table 11, pointing to class imbalance as less than $<$ 10% samples are assigned 2 as the median LJ rating.
Figure 10: Topical chat criteria – Natural. The human median is 3 for 100% of the samples. While here approximately 60% of the samples are also labelled 3 by LJ, the k-$\alpha$ correlation $\overline{\text{H}}\overline{\text{M}}$ is substantially low at -0.20 as shown Table 11.
Figure 11: Topical chat criteria – Overall. Here the human labels for bin $\overline{\text{H}}$=4 vary, where only 40% of the labels match the median of 4, hence humans are uncertain about this label assignment. Also note, this bin $\overline{\text{H}}$=4 applies to 8.3% (5/60) of the samples. For the rest of 91% of the samples, human uncertainty is low, but LJ median labels $\overline{\text{M}}$ vary are quite different to human median $\overline{\text{H}}=5$.